Circular No. 7985

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

ϵ ERIDANI, υ ANDROMEDAE, 51 PEGASI, HD 209458, HD 196050

L. Greenhill, Smithsonian Astrophysical Observatory, reports the nondetection of water-maser emission from five extrasolar planetary systems, including two from which water emission has been reported [Muir 2002, New Scientist 175(2361), 22], noting that young planetary systems may contain water in planets’ atmospheres, interplanetary material, and comets. Cometary water-maser emission is not commonplace in the solar system (Graham et al. 2000, A.J. 119, 2465), but comet-planet impacts may be a source of transient maser emission (Cosmovici 2000, ASP Conf. Ser. 213, 151). Observations were made of ϵ Eri on Sept. 21.7, 27.7, and 28.7 UT by P. Kondratko and J. Lovell with the 70-m NASA Deep Space Network (DSN) antenna (+ 400-MHz Smithsonian and 16-MHz Australia Telescope correlation spectrometers) near Canberra. On Sept. 21 and 27, the 400-MHz correlator generated spectra with 1σ-noise levels of 14 and 28 mJy (flux densities corrected for sky opacity), respectively, in each 1.3-km/s channel between heliocentric velocities ∼ −2650 and +2650 km/s. On Sept. 27 and 28, the 16-MHz correlator generated spectra with 1σ-noise levels in each 0.42-km/s channel of 46 mJy over −90 to +126 km/s and of 30 mJy over −25 to +191 km/s, respectively. J. F. Gómez, I. de Gregorio-Monsalvo, and T. Kuiper conducted observations of υ And using the DSN 70-m antenna (+ 10-MHz spectrometer) near Madrid on Sept. 24.0 and 25.9. The 1σ-noise levels for 0.53-km/s channels were 29, 74, and 47 mJy over velocity intervals of −253 to −132, −122 to +6, and +21 to +140 km/s, respectively. On Sept. 27.7, Lovell and Kondratko also observed υ And in Canberra with both correlators and obtained 1σ noise levels of 73 mJy over ∼ −2650 to +2650 km/s (with 1.3-km/s channels) and 64 mJy over −42 to +174 km/s (with 0.42-km/s channels). Systemic velocities observed are +15.5 km/s for ϵ Eri and −28.3 km/s for υ And. Between Sept. 21.3 and 21.6, Kondratko and Lovell also obtained nondetections at Canberra over ∼ −2650 to +2650 km/s (with 1.3-km/s channels) for 51 Peg (25 mJy), HD 209458 (22 mJy), and HD 196050 (28 mJy).

SUPERNOVA 2002gb Corrigendum. On IAUC 7983, line 8, the offset for SN 2002gb should

read 1′′.4 N

2002 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green